# Two- versus three-body approach to femtoscopic hadron-deuteron correlations

### v1

@@ 2 Two-body approach
Further on we assume the Gaussian form of the single-particle source function
@@ 5 Confrontation with experiment
The $p$-$d$ and $d$-$d$ correlation functions have been recently measured in Au-Au collisions at $\sqrt{s_{NN}}=3$ GeV. The radius $R_{pd}$ is indeed bigger than $R_{dd}$ but the ratio is about 1.7 not 1.2. The collision energy is rather low and deuterons are fragments of incoming nuclei. Then, not all deuterons are formed due to final state interactions and the relations (27) are of limited applicability. In any case it would be desirable to test the relations systematically.
@@ 6 Conclusions
The three-body approach to hadron-deuteron correlations changes into the two-body approach if the three-particle hadron-deuteron wave function factorizes into the deuteron wave-function and the wave function of a hadron motion relative to the deuteron. The assumption of factorization is less restrictive for the kaon-deuteron than the proton-deutereon system, as there is no ambiguity which particle should be bound to the neutron. Presumably for this reason the kaon-deuteron correlations in proton-proton collisions at the LHC are successfully described within the two-body approach while the description of the proton-deuteron correlations require the three-body approach.

### v2

@@ 2 Two-body approach
Further on we assume that the single-particle source function is Gaussian
@@ 3 Three-body approach
It is worth noting that the three-body approach to hadron-deuteron correlations is very similar to the approach to the correlations of three-particles in a scattering state that have recently been measured in case of $p$-$p$-$p$, $p$-$p$-$\bar{p}$, $p$-$p$-$\Lambda$ [15], $p$-$p$-$K$ [16] and $p$-$p$-$\pi$ [17] systems. The starting point for calculating the three-particle correlation function is a formula close to Eq. (16), which after eliminating the center-of-mass motion becomes analogous to Eq. (19), except that the wave function does not describe the motion of the hadron relative to the deuteron, but the relative motion of the three particles. The source function (20) can be directly used in the calculation of the $N$-$N$-$h$ correlation function. However, the studies [18, 19], where the $N$-$N$-$N$ and $N$-$N$-$\Lambda$ correlation functions have been calculated, clearly show that the real challenge is not to find the source function, but the three-particle wave function that takes into account the effect of genuine three-particle forces.
@@ 5 Confrontation with experiment
The $p$-$d$ and $d$-$d$ correlation functions have been recently measured in Au-Au collisions at $\sqrt{s_{NN}}=3$ GeV [10]. The radius $R_{pd}$ is indeed bigger than $R_{dd}$ but the ratio is about 1.7 not 1.2. The collision energy is rather low and deuterons are fragments of incoming nuclei. Then, not all deuterons are formed due to final state interactions and the relations (27) are of limited applicability. In any case it would be desirable to test the relations systematically.
@@ 6 Conclusions
The three-body approach to hadron-deuteron correlations changes into the two-body approach if the three-particle hadron-deuteron wave function factorizes into the deuteron wave-function and the wave function of a hadron motion relative to the deuteron. The assumption of factorization is less restrictive for the kaon-deuteron than the proton-deutereon system, as there is no ambiguity which particle should be bound to the neutron. Presumably for this reason the kaon-deuteron correlations in proton-proton collisions at the LHC are successfully described within the two-body approach while the description of the proton-deuteron correlations requires the three-body approach.